OpenStudy (anonymous):

8 15 --- + ---- = -1 x x^2

OpenStudy (anonymous):

answer choices A. x=5,-5 B. x=-5,-3 C. x=-5,3 D. x=5,3

OpenStudy (campbell_st):

multiply every term by x^2 and you'll have 8x + 15 = -x^2 or x^2 + 8x + 15 = 0 just factorise the quadratic for your solution.

OpenStudy (anonymous):

x=5,-5

OpenStudy (anonymous):

a. is supposed to be 5,-3 not 5

OpenStudy (campbell_st):

ok... can you factor the quadratic equation...?

OpenStudy (anonymous):

A. x=5,-3 B. x=-5,-3 C. x=-5,3 D. x=5,3

OpenStudy (anonymous):

not really

OpenStudy (campbell_st):

ok... can you tell be the factors of 15 that add to 8..?

OpenStudy (anonymous):

5 and 3

OpenStudy (campbell_st):

great so the quadratic when factorised is (x + 5)(x + 3) = 0 just solve these 2 equations x + 5 = 0 x + 3 = 0 for your answer

OpenStudy (anonymous):

-5, -3 ?

OpenStudy (campbell_st):

thats it... well done